# Proven Approximation Guarantees in Multi-Objective Optimization: SPEA2 Beats NSGA-II

### v1

@@ 6 Conclusion
We showed rigorously that the steady-state variant of the SPEA2, which generates one offspring each iteration, efficiently computes a desired optimal spread of individuals on the Pareto front of the classic OneMinMax benchmark when the population size is smaller than the size of the Pareto front (Theorem 1). Moreover, we proved in the same setting that the steady-state variant of the NSGA-II can be initialized such that, with overwhelming probability, it does not compute an optimal spread within polynomial time (Theorem 11). This difference in runtime performance is due to the $\sigma$-distances used in the SPEA2 for choosing which individual to remove, in contrast to the crowding distance used by the NSGA-II.

### v3

@@ 6 Conclusion
We showed rigorously that the steady-state variant of the SPEA2, which generates one offspring each iteration, efficiently computes a desired optimal spread of individuals on the Pareto front of the classic OneMinMax benchmark when the population size is smaller than the size of the Pareto front (Theorem 1). Moreover, we proved in the same setting that the steady-state variant of the NSGA-II can be initialized such that, with overwhelming probability, it does not compute an optimal spread within polynomial time (Theorem 11). This difference in runtime performance is due to the $\sigma$-distances used in the SPEA2 for choosing which individual to remove, in contrast to the crowding distance used by the NSGA-II, noting that the $\sigma$-distances are computationally more expensive to compute than the crowding distance.